OpenStudy (anonymous):

a fair dice is being rolled.what is the probability that a six appears in at most in at most n attempts?

OpenStudy (king):

hmmm. is it 1/6?

OpenStudy (anonymous):

once in 1 attempt = 1/6 twice in two attempts = 1/6 * 1/6 thrice in three attempts = 1/6 *1/6 *1/6 n in n attempts = (1/6)^n

OpenStudy (king):

oh ok!

OpenStudy (anonymous):

okay so it gets magnified according to the attempts